# MEDVOC: Vocabulary Adaptation for Fine-tuning Pre-trained Language Models on Medical Text Summarization

### v1

@@ Abstract
This work presents a dynamic vocabulary adaptation strategy, MEDVOC [^1] Accepted for publication at IJCAI 2024 (Main Track). This is the author’s version of the work. It is posted here for your personal use. Not for redistribution. , for fine-tuning pre-trained language models (PLMs) like BertSumAbs, BART, and PEGASUS for improved medical text summarization. In contrast to existing domain adaptation approaches in summarization, MEDVOC treats vocabulary as an optimizable parameter and optimizes the PLM vocabulary based on fragment score conditioned only on the downstream task’s reference summaries. Unlike previous works on vocabulary adaptation (limited only to classification tasks), optimizing vocabulary based on summarization tasks requires an extremely costly intermediate fine-tuning step on large summarization datasets. To that end, our novel fragment score-based hyperparameter search very significantly reduces this fine-tuning time — from $450$ days to less than $2$ days on average. Furthermore, while previous works on vocabulary adaptation are often primarily tied to single PLMs, MEDVOC is designed to be deployable across multiple PLMs (with varying model vocabulary sizes, pre-training objectives, and model sizes) — bridging the limited vocabulary overlap between the biomedical literature domain and PLMs. MEDVOC outperforms baselines by $15.74\%$ in terms of Rouge-L in zero-shot setting and shows gains of $17.29\%$ in high Out-Of-Vocabulary (OOV) concentrations. Our human evaluation shows MEDVOC generates more faithful medical summaries ($88\%$ compared to $59\%$ in baselines). We make the codebase publicly available at https://github.com/gb-kgp/MEDVOC.
§§footnotetext: Equal contribution.
@@ 4.4 Training Details
We obtain the biomedical paper abstracts from the official dump [^2] https://ftp.ncbi.nlm.nih.gov/pubmed/updatefiles of PubMed dated 2020. It comprises $450K$ data points (PubMed abstract and title). We refer to it as the PubMed Abstracts Collection (PAC) in the paper. We remove data points that overlap with downstream datasets (as EBM and BioASQ contain PubMed abstracts as the source document), as a decontamination step to prevent memorization issues Radford et al. (2019). We randomly select $312K$ data points to form the final dataset. We keep the dataset size of PAC similar to that of CNN/DailyMail for a fair comparison. Appendix A.5 provides further implementation details and the optimal hyperparameters and vocabulary size details.
@@ A.1 Pre-trained Language Models
  • BertSumAbs (BSA) Liu and Lapata [2019]: It uses a standard encoder-decoder framework where Bert acts as an encoder, and the decoder is six-layered transformer architecture initialized randomly [^3] https://github.com/nlpyang/PreSumm. BertSumAbs has $180$ Million parameters, uses the Word-Piece tokenizer and its pretraining objective is a combination of Masked Language Modeling and Next Sentence Prediction. The vocabulary size of this PLM ($|V_{\text{PLM}}|$) is $30522$.
  • BART Lewis et al. [2020]: BART is a denoising autoencoder, implemented as a sequence-to-sequence model with a bidirectional encoder over corrupted text and a left-to-right auto-regressive decoder to generate the original document it was derived from. We use the BART-LARGE [^4] https://huggingface.co/facebook/bart-large model available from the huggingface library. BART has $406$ Million parameters, uses Byte-Pair Encoding tokenization, and its pretraining objective is a combination of Text Infilling and Sentence Shuffling. The vocabulary size of this PLM ($|V_{\text{PLM}}|$) is $50265$.
  • PEGASUS Zhang et al. [2020a]: PEGASUS masks multiple whole sentences that are principle to the document and only generate the masked sentences as a single output sequence. We use the PEGASUS-LARGE [^5] https://huggingface.co/google/pegasus-large model available from the huggingface library. PEGASUS has $568$ Million parameters, uses Sentencepiece-Unigram tokenization, and its pretraining objective is Gap Sentence Generation. The vocabulary size of this PLM ($|V_{\text{PLM}}|$) is $96103$.
@@ A.2.1 Target Task Dataset Details
  • CHQSum Yadav et al. [2022]. CHQSum consists of $1507$ domain-expert annotated question-summary pairs from the Yahoo community question answering forum [^6] https://webscope.sandbox.yahoo.com/catalog.php?datatype=l&did=11 which provides community question answering threads containing users’ questions on multiple diverse topics and the answers submitted by other users. The authors with the help of $6$ domain experts identified valid medical question from the forum and asked the experts to formulate an abstractive summary for the questions.
@@ A.3 Evaluation Metrics
We first describe the implementation details for computing Rouge scores discussed in Section 4.2, where we use the official Rouge Lin [2004] script [^7] https://github.com/bheinzerling/pyrouge/tree/master. The following parameters: -c 95 -2 -1 -U -r 1000 -n 4 -w 1.2 -a, are used and we report the median at a $95$% confidence interval.
@@ A.5.2 Training Hyperparameters
All the experiments are run on three $32$ GB Tesla V100 GPUs. We use the training scripts provided by BertSumAbs in their codebase [^8] https://github.com/nlpyang/PreSumm. We use the standard fine-tuning summarization scripts for BART and PEGASUS provided in huggingface codebase [^9] https://github.com/huggingface/transformers/blob/main/examples/pytorch/summarization/run_summarization.py. The final training times for IFT-PAC for MEDVOC is mentioned in Table 6. We describe two types of hyperparameters.
@@ B.2 Human Evaluation
The annotations were conducted on the Prolific [^10] https://www.prolific.com/ platform using $30$ participants in total. Each participant was shown $10$ random samples from a pool of $100$ summary pairs (order of summary randomized and anonymized) and was given $30$ minutes to complete the study. We also collected demographic information and annotation experience feedback from the participants. The study instrument was designed using the Potato tool  [^11] https://github.com/davidjurgens/potato. The participants were compensated at the rate of $9$ pounds/hour, which according to platform guidelines was fair compensation. Proper consent notice was shown to the participants and no personal information (other than age) was collected in the demographic information. The filtering criteria for participants were kept as follows:

### v2

@@ Abstract
This work presents a dynamic vocabulary adaptation strategy, MEDVOC, for fine-tuning pre-trained language models (PLMs) like BertSumAbs, BART, and PEGASUS for improved medical text summarization. In contrast to existing domain adaptation approaches in summarization, MEDVOC treats vocabulary as an optimizable parameter and optimizes the PLM vocabulary based on fragment score conditioned only on the downstream task’s reference summaries. Unlike previous works on vocabulary adaptation (limited only to classification tasks), optimizing vocabulary based on summarization tasks requires an extremely costly intermediate fine-tuning step on large summarization datasets. To that end, our novel fragment score-based hyperparameter search very significantly reduces this fine-tuning time—from $450$ days to less than $2$ days on average. Furthermore, while previous works on vocabulary adaptation are often primarily tied to single PLMs, MEDVOC is designed to be deployable across multiple PLMs (with varying model vocabulary sizes, pre-training objectives, and model sizes) —bridging the limited vocabulary overlap between the biomedical literature domain and PLMs. MEDVOC outperforms baselines by $15.74\%$ in terms of Rouge-L in zero-shot setting and shows gains of $17.29\%$ in high Out-Of-Vocabulary (OOV) concentrations. Our human evaluation shows MEDVOC generates more faithful medical summaries ($88\%$ compared to $59\%$ in baselines).
§§footnotetext: Equal contribution.footnotetext: This is the author’s version of the work. It is posted here for your personal use. Not for redistribution. The definitive Version of Record was published in Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence, (IJCAI-24), https://doi.org/10.24963/ijcai.2024/683
@@ 4.4 Training Details
We obtain the biomedical paper abstracts from the official dump§§§https://ftp.ncbi.nlm.nih.gov/pubmed/updatefiles of PubMed dated 2020. It comprises $450K$ data points (PubMed abstract and title). We refer to it as the PubMed Abstracts Collection (PAC) in the paper. We remove data points that overlap with downstream datasets (as EBM and BioASQ contain PubMed abstracts as the source document), as a decontamination step to prevent memorization issues Radford et al. (2019). We randomly select $312K$ data points to form the final dataset. We keep the dataset size of PAC similar to that of CNN/DailyMail for a fair comparison. Appendix A.5 provides further implementation details and the optimal hyperparameters and vocabulary size details.
@@ A.1 Pre-trained Language Models
  • BertSumAbs (BSA) Liu and Lapata [2019]: It uses a standard encoder-decoder framework where Bert acts as an encoder, and the decoder is six-layered transformer architecture initialized randomly§§§https://github.com/nlpyang/PreSumm. BertSumAbs has $180$ Million parameters, uses the Word-Piece tokenizer and its pretraining objective is a combination of Masked Language Modeling and Next Sentence Prediction. The vocabulary size of this PLM ($|V_{\text{PLM}}|$) is $30522$.
  • BART Lewis et al. [2020]: BART is a denoising autoencoder, implemented as a sequence-to-sequence model with a bidirectional encoder over corrupted text and a left-to-right auto-regressive decoder to generate the original document it was derived from. We use the BART-LARGE§§§https://huggingface.co/facebook/bart-large model available from the huggingface library. BART has $406$ Million parameters, uses Byte-Pair Encoding tokenization, and its pretraining objective is a combination of Text Infilling and Sentence Shuffling. The vocabulary size of this PLM ($|V_{\text{PLM}}|$) is $50265$.
  • PEGASUS Zhang et al. [2020a]: PEGASUS masks multiple whole sentences that are principle to the document and only generate the masked sentences as a single output sequence. We use the PEGASUS-LARGE§§§https://huggingface.co/google/pegasus-large model available from the huggingface library. PEGASUS has $568$ Million parameters, uses Sentencepiece-Unigram tokenization, and its pretraining objective is Gap Sentence Generation. The vocabulary size of this PLM ($|V_{\text{PLM}}|$) is $96103$.
@@ A.2.1 Target Task Dataset Details
  • CHQSum Yadav et al. [2022]. CHQSum consists of $1507$ domain-expert annotated question-summary pairs from the Yahoo community question answering forum§§§https://webscope.sandbox.yahoo.com/catalog.php?datatype=l&did=11 which provides community question answering threads containing users’ questions on multiple diverse topics and the answers submitted by other users. The authors with the help of $6$ domain experts identified valid medical question from the forum and asked the experts to formulate an abstractive summary for the questions.
@@ A.3 Evaluation Metrics
We first describe the implementation details for computing Rouge scores discussed in Section 4.2, where we use the official Rouge Lin [2004] script§§§https://github.com/bheinzerling/pyrouge/tree/master. The following parameters: -c 95 -2 -1 -U -r 1000 -n 4 -w 1.2 -a, are used and we report the median at a $95$% confidence interval.
@@ A.5.2 Training Hyperparameters
All the experiments are run on three $32$ GB Tesla V100 GPUs. We use the training scripts provided by BertSumAbs in their codebase§§§https://github.com/nlpyang/PreSumm. We use the standard fine-tuning summarization scripts for BART and PEGASUS provided in huggingface codebase§§§https://github.com/huggingface/transformers/blob/main/examples/pytorch/summarization/run_summarization.py. The final training times for IFT-PAC for MEDVOC is mentioned in Table 6. We describe two types of hyperparameters.
@@ B.2 Human Evaluation
The annotations were conducted on the Prolific§§§https://www.prolific.com/ platform using $30$ participants in total. Each participant was shown $10$ random samples from a pool of $100$ summary pairs (order of summary randomized and anonymized) and was given $30$ minutes to complete the study. We also collected demographic information and annotation experience feedback from the participants. The study instrument was designed using the Potato tool §§§https://github.com/davidjurgens/potato. The participants were compensated at the rate of $9$ pounds/hour, which according to platform guidelines was fair compensation. Proper consent notice was shown to the participants and no personal information (other than age) was collected in the demographic information. The filtering criteria for participants were kept as follows: